REU: TENTATIVE SCHEDULE FOR WEEKS SEVEN AND EIGHT

Date  SESAME  Babai  Lewicka  May
Monday, August 4  9–12, 1–4 4:15 – 5:30  12:10 – 12:50
Tuesday, August 5  9–12, 1–4  4:15 – 5:30 12:10 – 12:50
Wednesday, August 6 9–12, 1–4 4:15 – 5:30  12:10 – 12:50
Thursday, August 7 9–12, 1–4  4:15 – 5:30 12:10 – 12:50
Friday, August 8  9–12, 1–4 4:15 – 5:30  12:10 – 12:50
Date  Babai  Lewicka Student presentations
Monday, August 10  9:00 – 12:00 1:30 – 3:00  3:30 on
Tuesday, August 11  9:00 – 12:00 1:30 – 3:00  3:30 on
Wednesday, August 12 9:00 – 12:00 1:30 – 3:00  3:30 on
Thursday, August 13  All day
Friday, August 14  All day, party at 5:00

I urge graduate student counsellors to offer daily afternoon problem sessions, between 1:00 and 4:00, for the YSP participants (and SESAME participants who work in the mornings) on any and all of the courses to date. All undergraduate participants should feel free to e-mail graduate students to set up one-on-one or small group meetings, and I remind undergraduates that we have graduate student volunteers willing to mentor directed reading programs between the end of the REU and the beginning of the fall quarter.

Here is the schedule in summary form:

Seventh week:

SESAME 9–12, 1–4 M – F Babai 4:15 – 5:30 MWF Lewicka 4:15 – 5:30 TTh May 12:10–12:50 M – F

Eighth week: Babai 9:00 – 12:00 MTW Lewicka 1:30 – 3:00 MTW Student presentations to be arranged M – F Party 5:00 – ?:?? Friday